# DynoSAM: Open-Source Smoothing and Mapping Framework for Dynamic SLAM

### v2

@@ IV-A Rigid-Body Motion
Understanding rigid-body motion is fundamental to Dynamic SLAM because it provides a mathematical framework to model the motion of objects in the environment. Our framework utilizes this understanding to estimate the $\mathrm{SE}(3)$ motion of objects directly using frame-to-frame observations of 3D points on rigid bodies.
A motion is defined as any change in pose, following [23], which describes the conversion of one pose into another. Both poses and motions are represented using $\mathrm{SE}(3)$ homogeneous transformations. However, while a pose defines a body’s position and orientation relative to a reference frame, a motion maps one pose onto another. Our work leverages this concept and employs the three-index notation [23] for homogeneous transformations, allowing both the frame of reference and the frame in motion to be specified. For example, ${}^{O}_{A}\mathbf{H}_{B}\in\mathrm{SE}(3)$ is the transformation which describes the motion from frame $\{A\}$ to frame $\{B\}$ from the perspective of frame $\{O\}$.
@@
A rigid body, by definition, maintains its shape and size throughout motion and therefore, any rigid-body motion must ensure that the distance between any two points on the object remains constant. The work of [23] shows that while we can represent this rigid-body motion in any arbitrary frame, the observed motion of each point on the rigid-body depends on the frame of the observer. For instance, a point on a rotating wheel might appear stationary to an observer positioned locally on the same wheel but will appear to be moving rapidly to a stationary observer that is external to the wheel. Furthermore, we can also envisage an observer moving relative to the spinning wheel! This example illustrates the difference between local motion, which describes the motion as perceived by (or on) the object, and observed motion, which describes the motion as perceived by an external observer.
Since it is our goal to estimate the object’s rigid-body motion from observing a small set of individual points, consider attempting to estimate the rigid-bodies local motion: each point observed will exhibit their own unique local motion, thus making it complex to directly infer the overall motion of the rigid-body. This challenge can be overcome by instead considering the motion of each point on the rigid body as observed by an external reference frame. By representing points on the rigid-body in a common, external reference frame, the motion of all points, and hence, the body itself, can be described using a single $\mathrm{SE}(3)$ transformation. This approach is highly relevant in Dynamic SLAM as only partial observations of an object are often possible and the observed motion representation allows all point observations to directly contribute to estimating the object’s overall motion.
Figure 2: Illustrative example showing the application of local (red) and observed (orange) motions on a rigid-body. Two arbitrary frames $A$ and $B$ are shown before and after a motion is applied to body. This figure illustrates that both local and observed motions can be used to correctly propagate the fixed frames ($\{A_{1}\}$ to $\{A_{2}\}$ and $\{B_{1}\}$ to $\{B_{2}\}$) in accordance with rigid-body kinematics. However, local motion is dependent upon the position and orientation of the reference frame, leading to the invalid propagation of $\{B_{1}\}$ to ${{}^{O}\hat{\mathbf{B}}_{2}}$ if incorrectly applied.
To further illustrate these concepts, consider the two arbitrary reference frames $\{A_{1}\}$ and $\{B_{1}\}$ as depicted in Fig. 2. Both frames are rigidly attached to the shown rigid body. After some motion is applied to the body, their corresponding frames are shown as $\{A_{2}\}$ and $\{B_{2}\}$. All frames are defined with their associated poses ${{}^{O}\mathbf{A}_{1}}$, ${{}^{O}\mathbf{B}_{1}}$, ${{}^{O}\mathbf{A}_{2}}$ and ${{}^{O}\mathbf{B}_{2}}$ with respect to the fixed origin frame $\{O\}$. Vitally, since the body is rigid, the relative pose transformation between frames $A$ and $B$ remains constant over time and defines the kinematic constraints of a rigid body. We denote these pose transformations as $\prescript{B_{1}}{B_{1}}{\mathbf{C}}_{A_{1}}=\mathbf{C}=\prescript{B_{2}}{B_{2}}{\mathbf{C}}_{A_{2}}$. Contrastingly, we use $\prescript{}{}{\mathbf{H}}$ to specify any motion described by a homogeneous transformation.
IV-A1 Local Motion
Local motion describes the process of moving a coordinate frame from one pose relative to another and can be thought of as the motion of the frame as perceived by the object. A transformation of this form can be clearly specified using the three-indexed notation ${}^{A}_{A}\mathbf{H}_{B}$. As noted in [23], the robotics community commonly uses this representation by default, allowing for the simpler notation such as ${}^{A}\mathbf{H}_{B}$.
We use Fig. 2 to illustrate how local motion (shown on the figure in red) can be used to correctly propagates frames $\{A_{1}\}$ to $\{A_{2}\}$ and $\{B_{1}\}$ to $\{B_{2}\}$:
@@ IV-A1 Local Motion
In this context ‘correctly’ means that no rigid-body kinematic constraint are violated. However, the applicability of local motion is constrained by the position and orientation of the reference frame, making it dependent on the frame’s origin. Such limitation can lead to inconsistencies when propagating frames with different origins; this specific case is illustrated in Fig. 2 where $\prescript{A_{1}}{A_{1}}{\mathbf{H}}_{A_{2}}$ is used to propagate frame $\{B_{1}\}$. The resulting pose ${{}^{O}\hat{\mathbf{B}}_{2}}$ clearly violates rigid-body constraints.
IV-A2 Observed Motion
Observed motion describes the motion of a rigid-body as seen by an external observer. Fig. 2 uses orange arrows to depict the propagation $\{A_{1}\}$ and $\{B_{1}\}$ using the observed motion $\prescript{O}{1}{\mathbf{H}}_{2}$, which is defined with respect to the origin frame $\{O\}$. In contrast to local motion in Eq. (1), the same observed motion $\prescript{O}{1}{\mathbf{H}}_{2}$ can propagate both $\{A_{1}\}$ and $\{B_{1}\}$ in accordance with rigid-body kinematics. This is possible because the observed motion representation is independent of the position and orientation of the frame that it is acting upon.
@@ IV-A2 Observed Motion
While these underlying principles have been previously established, notably in [23] and subsequently applied to the estimation problems in our prior works [11, 15, 18], a full derivation of observed motion within the context of Dynamic SLAM has not been previously presented. We believe a more complete derivation is crucial for understanding the benefits of this representation. In this section, we exploit the constraints of a rigid body to show how $\prescript{O}{1}{\mathbf{H}}_{2}$ can be derived given the relations defined in Eq. (1) and the constant relative transformation $\mathbf{C}$ between $\{A\}$ and $\{B\}$. By applying a frame change of a pose transformation [23] we can express each local motion in a common origin frame:
$\prescript{O}{1}{\mathbf{H}}_{2}\coloneqq$ ${{}^{O}\mathbf{A_{1}}}\>\prescript{A_{1}}{A_{1}}{\mathbf{H}}_{A_{2}}\>{{}^{O}\mathbf{A_{1}}^{-1}}$ (2)
$=$ ${{}^{O}\mathbf{B_{1}}}\>\prescript{B_{1}}{B_{1}}{\mathbf{C}}_{A_{1}}\>\prescript{A_{1}}{A_{1}}{\mathbf{H}}_{A_{2}}\>\prescript{B_{1}}{B_{1}}{\mathbf{C}}_{A_{1}}^{-1}\ {{}^{O}\mathbf{A_{1}}^{-1}}$
$=$ ${{}^{O}\mathbf{B_{1}}}\>\prescript{B_{1}}{B_{1}}{\mathbf{C}}_{A_{1}}\>\prescript{A_{1}}{A_{1}}{\mathbf{H}}_{A_{2}}\>\prescript{B_{2}}{B_{2}}{\mathbf{C}}_{A_{2}}^{-1}\ {{}^{O}\mathbf{A_{1}}^{-1}}$
$=$ ${{}^{O}\mathbf{B_{1}}}\>\prescript{B_{1}}{B_{1}}{\mathbf{H}}_{B_{2}}\>{{}^{O}\mathbf{B_{1}}^{-1}}\text{.}$
@@
Eq. (2) shows that $\prescript{O}{1}{\mathbf{H}}_{2}$ remains constant regardless of the choice of $\mathbf{A}$ and $\mathbf{B}$ and, therefore, this representation of motion is independent of the rigid body’s frame of reference. Furthermore, using Eq. (1), we can rewrite Eq. (2) as:
@@
We highlight that the composition order of the observed motion in Eq. (4) is not the same as that for the local motion in Eq. (1) due to their fundamental differences, and we direct the interested reader to [23] for more details.
IV-B Notations
We define the notations for the Dynamic SLAM problem by considering a scene at a discrete time-step $k$ with a set of objects and their associated poses with respect to a fixed world frame $\{W\}$ denoted as ${{}^{W}\mathcal{L}_{k}}$:
${{}^{W}\mathcal{L}_{k}}=\{\prescript{W}{}{\mathbf{L}}_{k}^{j}\in\mathrm{SE}(3),j\in\mathcal{J}_{k},k\in\mathcal{K}\}$
where $\mathcal{J}_{k}$ is the set of object indices observed at time-step $k$ and $\mathcal{K}$ contains all time-steps of the sequence. Each object pose $\prescript{W}{}{\mathbf{L}}_{k}^{j}$ is associated with a body-fixed reference frame $\{L_{k}\}$. For the $j$-th object, we define its observation history ${{}^{W}\mathcal{L}}$ using the $n+1$ consecutive frames that the object has been observed in ${{}^{W}\mathcal{L}^{j}}=\{\prescript{W}{}{\mathbf{L}}_{k}^{j}\}_{k=s...s+n}$, where $s$ is the first time-step at which this object is observed. Using this notation we can write the first pose of the object trajectory as $\prescript{W}{}{\mathbf{L}}_{s}^{j}$. Similarly, $\prescript{W}{}{\mathbf{X}}_{k}\in\mathrm{SE}(3)$ is the camera pose with respect to the world frame $\{W\}$ at time-step $k$, and is associated with a camera frame $\{X_{k}\}$. The set of all camera poses is denoted as ${}^{W}\mathcal{X}=\{\prescript{W}{}{\mathbf{X}}_{k}\in\mathrm{SE}(3),k\in\mathcal{K}\}$
@@ IV-B Notations
We use $\prescript{}{}{\mathbf{m}}$ to refer to both static and dynamic points. In the case of a static point, $\prescript{W}{}{\mathbf{m}}_{k}\doteq\prescript{W}{}{\mathbf{m}}$ as we omit the time-step $k$ to indicate the variable is time invariant. Finally, ${{}^{W}\mathcal{M}_{k}}$ refers to the set of all points in the world frame at $k$.
We denote motion of the camera as $\prescript{}{}{\mathbf{T}}$ and the motion of any object $j$ as $\prescript{}{}{\mathbf{H}}$. From now onwards we use $\prescript{}{}{\mathbf{H}}$ to only refer to the motion of objects, rather denoting general motion as in Sec. IV-A. Between time-steps $k-1$ and $k$ the local motions of the camera and object are $\prescript{X_{k-1}}{X_{k-1}}{\mathbf{T}}_{X_{k}}$ and $\prescript{L_{k-1}}{L_{k-1}}{\mathbf{H}}_{L_{k}}$ respectively. In the case where our notation is already explicit in denoting which pose the transformation is referring to, i.e., camera pose with $\prescript{}{}{\mathbf{T}}$ or object pose with $\prescript{}{}{\mathbf{H}}$, we simplify the notation by only using time-steps:
$\prescript{X_{k-1}}{k-1}{\mathbf{T}}_{k}\doteq\prescript{X_{k-1}}{X_{k-1}}{\mathbf{T}}_{X_{k}}$ (5)
$\prescript{L_{k-1}}{k-1}{\mathbf{H}}_{k}\doteq\prescript{L_{k-1}}{L_{k-1}}{\mathbf{H}}_{L_{k}}$
These transformations describe local motions and, following Eq. (1), can be used to propagate the camera and object poses from $k-1$ to $k$:
@@
For the purposes of object motion estimation, our framework uses an observed motion representation, as discussed in Sec. IV-A2. In the context of SLAM this reference frame can be any observable frame, e.g. the camera frame at the start of a sliding-window. In this work, we represent our object motions in the world frame $\{W\}$, as introduced by our previous works [11, 15, 18]. Such a representation is convenient for downstream applications like mapping and planning, as it allows all state variables in our formulations to be defined with respect to a common coordinate system. For any object $j$ this motion is denoted as $\prescript{W}{k-1}{\mathbf{H}}_{k}$. Due to its representation in the the world frame we refer specifically to it as the absolute motion.
Using Eq. (3) we define our absolute motion in terms of a change in object pose:
@@
Following the composition rules defined in Eq. (4), $\prescript{W}{k-1}{\mathbf{H}}_{k}$ will move $\prescript{W}{}{\mathbf{L}}_{k-1}$ to $\prescript{W}{}{\mathbf{L}}_{k}$ when applied on the left-hand side:
@@
$\prescript{W}{}{\mathbf{m}}_{k}^{i}=\prescript{W}{k-1}{\mathbf{H}}_{k}\>\prescript{W}{}{\mathbf{m}}_{k-1}^{i}\text{,}$ (10)
relating the motion of any and all points on an object with a single transformation. Since we can directly observe 3D object points, Eq. (10) allows us to easily derive cost functions in terms of all tracked points on the object. This equation underpins all our formulations and is fundamental to our Dynamic SLAM framework. A more detailed derivation of Eq. (10) is included in Appendix -A which highlights how this equation implicitly encodes the rigid-body kinematic constraints discussed in Sec. IV-A.
As before, we define the set of object motions at timestamp $k$ for all objects $j\in\mathcal{J}_{k}$ as:
$\prescript{W}{k-1}{\mathbf{\mathcal{H}}}_{k}=\{\prescript{W}{k-1}{\mathbf{H}}_{k}^{j},j\in\mathcal{J}_{k},k\in\mathcal{K}\}$
For all notation we often omit indices $i$ and $j$ when there is no ambiguity.
V Foundations for Dynamic SLAM
This section details the graph-based Dynamic SLAM formulations included within our framework. We demonstrate how the concept of observed motion, explained in Sec. IV, is applied to directly estimate both object motions (Sec. V-B) and poses (Sec. V-C). These formulations constitute the back-end of a visual Dynamic SLAM pipeline. As a result, we assume dynamic objects can be tracked and 3D point measurements of static and dynamic features can be provided by a visual SLAM front-end. While we outline our front-end (Sec. VI-B) and include it in our open-source implementation, the formulations discussed in our framework are agnostic to any particular front-end designs.

### v3

@@ III-A Rigid-Body Motion
Understanding rigid-body motion is fundamental to Dynamic SLAM as it provides a mathematical foundation to model the motion of objects in the environment. Our framework builds upon this understanding to estimate the motion of objects directly using frame-to-frame observations of 3D points on rigid bodies.
@@
A rigid body, by definition, maintains its shape and size throughout motion; and therefore the distance between any pair of points on the object will not change. However, while these points must move together rigidly, the apparent motion of each point depends on the location of the observer. For instance, a point on a rotating wheel may appear stationary to an onboard observer but moving to one to a stationary observer external to the wheel. This example highlights the distinction between what we term as local motion, which is perceived by a reference frame that is rigidly attached to the object, and the observed motion which is perceived by an external observer.
In accordance with rigid-body kinematics, a reference frame is represented by a homogeneous transformation and can be rigidly attached to a rigid body, or can be placed anywhere in the environment. A pose of a frame defines the $\mathrm{SE}(3)$ homogeneous transform that contains the frame’s position and orientation with respect to another reference frame, e.g. the observer frame.
To discuss these concepts more concretely, consider Fig. 2, which depicts a rigid body moving freely from time-step $1$ to $2$. Two arbitrary frames $\{A\}$ and $\{B\}$ are fixed to the object and observed by the fixed reference frame $\{O\}$. $^{O}\mathbf{A}_{1},^{O}\mathbf{B}_{1}\in\mathrm{SE}(3)$ denote the poses of frames $\{A_{1}\}$ and $\{B_{1}\}$ relative to $\{O\}$ prior to motion, while $^{O}\mathbf{A}_{2},^{O}\mathbf{B}_{2}\in\mathrm{SE}(3)$ denote the corresponding poses after motion, as shown by the solid black arrows in Fig. 2.
For any moving frame, we define its motion as a change in pose following Chirikjian et al[Chirikjian17idetc]. This motion is a $\mathrm{SE}(3)$ homogeneous transformation that converts one pose into another and thus describes the movement of a frame. Importantly, for a rigid body this motion can be represented in any arbitrary frame [Chirikjian17idetc]. To express this information unambiguously, we adopt the three-index notation [Chirikjian17idetc] for these motions which allows the observer frame to be explicitly identified in addition to the moving frame. For example, $\prescript{A_{1}}{A_{1}}{\mathbf{H}}_{A_{2}}$ and $\prescript{B_{1}}{B_{1}}{\mathbf{H}}_{B_{2}}$ are the local motions for frames $\{A\}$ and $\{B\}$ respectively, while the observed motion is $\prescript{O}{1}{\mathbf{H}}_{2}$. The remainder of the section details how each local motion is depends on its frame of origin, in contrast to the observed motion which is independent of the moving frame and can therefore describe the rigid-body motion of any point on the rigid-body.
Figure 2: Illustrative example of two arbitrary frames, $\{A\}$ and $\{B\}$, attached to a moving rigid body. Solid black arrows show the poses of each frame, while green and blue arrows show the corresponding local motions of each frame. The red arrow denotes the observed motion $\prescript{O}{1}{\mathbf{H}}_{2}$ which is a unique transform that describes the rigid-body motion of all points on the object. For this reason we draw it unattached to any specific reference frame. $\mathbf{C}$ is a constant transform that defines the rigid-body’s kinematic constraint. Observe that both local and observed motions can be used to correctly transform the pose of each frame in accordance the kinematic constraints. However, local motion is dependant on the chosen reference frame and incorrect use leads to the invalid propagation of $^{O}\mathbf{B}_{1}$ to $^{O}\hat{\mathbf{B}}_{2}$}
III-A1 Local Motion
Local motion describes the process of moving a reference frame from one pose relative to another as perceived by a frame on the moving object. Local motions are shown in Fig. 2 as green and blue arrows e.g. $\prescript{A_{1}}{A_{1}}{\mathbf{H}}_{A_{2}}\in\mathrm{SE}(3)$ is the transformation of pose $^{O}\mathbf{A}_{1}$ to $^{O}\mathbf{A}_{2}$ from the perspective of frame $\{A_{1}\}$ [^1] It is worth noting that usually in robotics $\prescript{A}{A}{\mathbf{H}}_{B}$ is considered, which allows for the simpler notation $\prescript{A}{}{\mathbf{H}}_{B}$. In this case, $\prescript{A}{}{\mathbf{H}}_{B}$ is the coordinate transform from $A$ to $B$ and whose matrix is the same as that of the motion $\prescript{A}{A}{\mathbf{H}}_{B}$..
These local motions form a standard kinematic chain which allows poses of frame $\{A\}$ and $\{B\}$ to be propagated such that no rigid-body kinematic constraint are violated, i.e. $\mathbf{C}$ remains constant:
@@ III-A1 Local Motion
Vitally, each local motion is dependent on the frame’s origin and can therefore break the rigid-body constraint if applied to a pose with a different origin; this specific case is illustrated in Fig. 2 where $\prescript{A_{1}}{A_{1}}{\mathbf{H}}_{A_{2}}$ is applied to $^{O}\mathbf{B}_{1}$. The resulting pose $^{O}\hat{\mathbf{B}}_{2}$} violates rigid-body constraints, as the relative transformation from $^{O}\hat{\mathbf{B}}_{2}$} to $^{O}\mathbf{A}_{2}$ is clearly not equal to $\mathbf{C}$.
III-A2 Observed Motion
Observed motion describes the motion of a rigid body as seen by an external observer. In Fig. 2, the observed motion from time-steps $1$ to $2$ is defined with respect to $\{O\}$ and is therefore written $\prescript{O}{1}{\mathbf{H}}_{2}$. Unlike local motion in (1), this motion is invariant to the choice of body-fixed reference frame and can there be applied to any pose associated with a fixed reference on the object, whilst upholding rigid-body constraints: $\prescript{O}{1}{\mathbf{H}}_{2}=\prescript{O}{A_{1}}{\mathbf{H}}_{A_{2}}=\prescript{O}{B_{1}}{\mathbf{H}}_{B_{2}}$. The observed motion is shown with a red arrow that, importantly, is not attached to any specific reference frame. This highlights that, unlike local motions, the observed motion does not have a geometric interpretation yet exists as a unique transform that can propagate the poses $^{O}\mathbf{A}_{1}$ and $^{O}\mathbf{B}_{1}$ to $^{O}\mathbf{A}_{2}$ and $^{O}\mathbf{B}_{2}$, as well as any other point that is rigidly attached to the object.
@@ III-A2 Observed Motion
While these underlying principles have been previously established, notably in [Chirikjian17idetc] and subsequently applied to the estimation problems in [Henein20icra, zhang2020vdoslam, morris2024icra], a full derivation of observed motion within the context of Dynamic SLAM has not been previously presented. We believe a more complete derivation is crucial for understanding the benefits of this representation and in this section we show how to derive $\prescript{O}{1}{\mathbf{H}}_{2}$ given the relations defined in (1) and by exploiting the rigid-body constraint $\mathbf{C}$ between $\{A\}$ and $\{B\}$. Each local motion can be represented with a single observed motion when expressed in a common origin frame:
$\prescript{O}{1}{\mathbf{H}}_{2}\coloneqq$ ${{}^{O}\mathbf{A_{1}}}\>\prescript{A_{1}}{A_{1}}{\mathbf{H}}_{A_{2}}\>{{}^{O}\mathbf{A_{1}}^{-1}}$ (2a)
$=$ ${{}^{O}\mathbf{B_{1}}}\>\prescript{B_{1}}{B_{1}}{\mathbf{C}}_{A_{1}}\>\prescript{A_{1}}{A_{1}}{\mathbf{H}}_{A_{2}}\>\prescript{B_{1}}{B_{1}}{\mathbf{C}}_{A_{1}}^{-1}\ {{}^{O}\mathbf{B_{1}}^{-1}}$ (2b)
$=$ ${{}^{O}\mathbf{B_{1}}}\>\prescript{B_{1}}{B_{1}}{\mathbf{C}}_{A_{1}}\>\prescript{A_{1}}{A_{1}}{\mathbf{H}}_{A_{2}}\>\prescript{B_{2}}{B_{2}}{\mathbf{C}}_{A_{2}}^{-1}\ {{}^{O}\mathbf{B_{1}}^{-1}}$ (2c)
$=$ ${{}^{O}\mathbf{B_{1}}}\>\prescript{B_{1}}{B_{1}}{\mathbf{H}}_{B_{2}}\>{{}^{O}\mathbf{B_{1}}^{-1}}\text{,}$ (2d)
where (2a) is the definition of a frame change of a pose transformation [Chirikjian17idetc], and because $\prescript{B}{}{\mathbf{C}}_{A}={{}^{O}\mathbf{B}^{-1}}\>{{}^{O}\mathbf{A}}$, we arrive at (2b). Given that frames $\{A\}$ and $\{B\}$ belong to the same rigid body, their relative transformation $\prescript{B}{}{\mathbf{C}}_{A}$, is invariant to time, i.e. $^{B_{1}}\mathbf{C}_{A_{1}}={{}^{B_{2}}\mathbf{C}_{A_{2}}}$, hence we can reach (2c), which can be re-written as (2d) following the kinematic chain visualised in Fig. 2.
@@
Equation (2) shows that $\prescript{O}{1}{\mathbf{H}}_{2}$ remains constant regardless of the choice of $\{A\}$ and $\{B\}$ and, therefore, this representation of motion is independent of the rigid body’s frame of reference. Furthermore, using (1), we can rewrite (2) as:
@@
We highlight that the composition order of the observed motion in (4) is not the same as that for the local motion in (1) due to their fundamental differences, and we direct the interested reader to [Chirikjian17idetc] for more details.
In a dynamic scene with many objects and significant occlusions, the distinction between local and observed motion play a crucial role. The camera can only observe the motions of a small set of points on each object in a busy scene and considering only local motion makes it difficult to infer the object’s overall motion as each point experiences a unique motion. However, if instead we expressing these points in a common external reference frame, the apparent motion of all points on the object can be observed directly and described using a single homogeneous transformation, i.e. the observed motion [Chirikjian17idetc], which is the foundation for our formulation of the Dynamic SLAM problem.
III-B Notations
We define the notations for the Dynamic SLAM problem by considering a dynamic scene comprised of camera poses $\mathcal{X}$ and object poses $\mathcal{L}$:
$\mathcal{X}=\{\prescript{W}{}{\mathbf{X}}_{k}\in\mathrm{SE}(3)\}_{{k\in\mathcal{K}}},\quad\mathcal{L}=\{\prescript{W}{}{\mathbf{L}}_{k}^{j}\in\mathrm{SE}(3)\}{\begin{subarray}{c}{j\in\mathcal{J}_{k}}\\ {k\in\mathcal{K}_{j}}\end{subarray}}$
where $\{W\}$ is the fixed world frame, $\mathcal{K}$ is the set of all time-steps and $\mathcal{J}$ is the set of all object indices. Throughout this work we liberally use the super-script $j$ or sub-script $k$ to specify the subset denoting to a particular object or at a specific time-step, e.g. $\mathcal{J}_{k}\in\mathcal{J}$ referring to all the observed objects at time-step $k$. The camera $\prescript{W}{}{\mathbf{X}}_{k}$ and object $\prescript{W}{}{\mathbf{L}}_{k}$ poses define the location of frames $\{X_{k}\}$ and $\{L_{k}\}$ with respect to $\{W\}$.
@@ III-B Notations
We use $\prescript{}{}{\mathbf{m}}\in\mathcal{M}$ to refer to both static and dynamic points. In the case of a static point, $\prescript{W}{}{\mathbf{m}}_{k}\doteq\prescript{W}{}{\mathbf{m}}$ as we omit the time-step $k$ to indicate the variable is time invariant. To minimize notation clutter, we liberally omit indices $i$ and $j$ when there is no ambiguity.
We denote motion of the camera as $\prescript{}{}{\mathbf{T}}$ and the motion of any object $j$ as $\prescript{}{}{\mathbf{H}}$. From now onwards we use $\prescript{}{}{\mathbf{H}}$ to only refer to the motion of objects, rather denoting general motion as in Sec. III-A. Between time-steps k1k-1 and $k$ the local motions of the camera and object are $\prescript{X_{k-1}}{X_{k-1}}{\mathbf{T}}_{X_{k}}$ and $\prescript{L_{k-1}}{L_{k-1}}{\mathbf{H}}_{L_{k}}$ respectively. In the case where our notation is already explicit in denoting which pose the transformation is referring to, i.e., camera pose with $\prescript{}{}{\mathbf{T}}$ or object pose with $\prescript{}{}{\mathbf{H}}$, we simplify the notation by only using time-steps:
$\prescript{X_{k-1}}{k-1}{\mathbf{T}}_{k}\doteq\prescript{X_{k-1}}{X_{k-1}}{\mathbf{T}}_{X_{k}},\quad\prescript{L_{k-1}}{k-1}{\mathbf{H}}_{k}\doteq\prescript{L_{k-1}}{L_{k-1}}{\mathbf{H}}_{L_{k}}$
These transformations describe local motions and, following (1), can be used to propagate the camera and object poses from k1k-1 to $k$:
@@
For the purposes of object motion estimation, our framework uses an observed motion representation, as discussed in Sec. III-A2. In the context of SLAM this reference frame can be any observable frame, e.g. the camera frame at the start of a sliding-window. In this work, we represent our object motions in the world frame $\{W\}$, following the works of [Henein20icra, zhang2020vdoslam, morris2024icra]. Such a representation is convenient for downstream applications like mapping and planning, as it allows all state variables in our formulations to be defined with respect to a common coordinate system. For any object $j$ this motion is denoted as $\prescript{W}{k-1}{\mathbf{H}}_{k}$. Due to its representation in the world frame we refer specifically to it as the absolute motion. We define the set of all object motions as:
$\prescript{W}{k-1}{\mathbf{\mathcal{H}}}_{k}=\{\prescript{W}{k-1}{\mathbf{H}}_{k}^{j}\}{\begin{subarray}{c}{j\in\mathcal{J}_{k}}\\ {k\in\mathcal{K}_{j}}\end{subarray}}$
Using (3) we define our absolute motion in terms of a change in object pose:
@@
Following the composition rules defined in (4), $\prescript{W}{k-1}{\mathbf{H}}_{k}$ will move $\prescript{W}{}{\mathbf{L}}_{k-1}$ to $\prescript{W}{}{\mathbf{L}}_{k}$ when applied on the left-hand side:
@@
$\prescript{W}{}{\mathbf{m}}_{k}^{i}=\prescript{W}{k-1}{\mathbf{H}}_{k}\>\prescript{W}{}{\mathbf{m}}_{k-1}^{i}\text{.}$ (9)
Thus, the transform $\prescript{W}{k-1}{\mathbf{H}}_{k}$ describes the motion of any and all points on an object from k1k-1 to $k$. Since we can directly observe 3D object points, (9) allows us to easily derive cost functions in terms of all tracked points on the object. This equation underpins all our formulations and is fundamental to our Dynamic SLAM framework. A more detailed derivation of (9) is included in Appendix -A which highlights how this equation implicitly encodes the rigid-body kinematic constraints discussed in Sec. III-A.
IV Foundations for Dynamic SLAM
This section details the graph-based Dynamic SLAM formulations included within our framework. We demonstrate how the concept of observed motion, explained in Sec. III, is applied to directly estimate both object motions (Sec. IV-B) and poses (Sec. IV-C). These formulations constitute the back-end of a visual Dynamic SLAM pipeline. As a result, we assume dynamic objects can be tracked and 3D point measurements of static and dynamic features can be provided by a visual SLAM front-end. While we outline our front-end (Sec. V-B) and include it in our open-source implementation, the formulations discussed in our framework are agnostic to any specific front-end implementation.